A girl was born 35years after her mother. What fraction of the mother's age wiill be the girl at age 15?
Answer by stanbon(75887)

A girl was born 35years after her mother. What fraction of the mother's age will be the girl at age 15?
--------------
Equations:
m = d + 35
----
When girl is 15, mom is 35+15 = 50
-----
Ans: girl = (15/50) = 3/10 of the mother's age.
---------------
Cheers,
Stan H.
-------------